OpenStudy (anonymous):

Find the solution to 7+8x+-1, 10+h/3=1, m/6+5=26, and -5x-2=13. Please help.

OpenStudy (anonymous):

@mthompson440 Please help me

OpenStudy (anonymous):

Do you know how to solve for X?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

Alright, I might not be able to help you that long because I am leaving soon. But I will try to help as much as I can.

OpenStudy (anonymous):

thats ok

OpenStudy (anonymous):

sorry, will help you later.

OpenStudy (lynfran):

is there a typo with the first problem..?? 7+8x+-1 ??

OpenStudy (anonymous):

no its really looks like -1=7+8x

OpenStudy (lynfran):

thats a better view...anyways... we need to solve for x so we can take 7 frm both sides of the equation we get -1-7=7-7+8x simplify we get -1-7=8x wats -1-7 equals to?

OpenStudy (anonymous):

-8

OpenStudy (lynfran):

good so -8=8x

OpenStudy (lynfran):

we now divide both sides by 8 to solve for x

OpenStudy (lynfran):

whats did u get for x?

OpenStudy (anonymous):

-1?

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

I kept on getting -8

OpenStudy (lynfran):

-8/8=-1

OpenStudy (lynfran):

-8=8x -8/8=8x/8 -1=x

OpenStudy (anonymous):

i get it

OpenStudy (lynfran):

ok

OpenStudy (anonymous):

how would you do 10+h/3=1

OpenStudy (lynfran):

first subtract 10 frm both sides of the equation so 10-10+h/3=1-10 simplify we get..?

OpenStudy (anonymous):

0+h/3=-9?

OpenStudy (lynfran):

yes h/3=-9

OpenStudy (lynfran):

now we multiply both sides of the equation by b ...that way we can get h alone on the left side of the equation

OpenStudy (lynfran):

by 3"

OpenStudy (lynfran):

multiply both sides of the equation by 3

OpenStudy (anonymous):

-9 times 3 equals -27

OpenStudy (lynfran):

yes thats what h= to

OpenStudy (anonymous):

-27?

OpenStudy (lynfran):

yep

OpenStudy (lynfran):

the next equation is done the same way like the previous 1 above m/6+5=26 whats the first step??

OpenStudy (anonymous):

subtract 5 from each side

OpenStudy (lynfran):

correct then what??

OpenStudy (anonymous):

multiply 6 to each side

OpenStudy (anonymous):

I got 126

OpenStudy (lynfran):

me too

OpenStudy (anonymous):

what about the next equation? the subtraction sign got me

OpenStudy (lynfran):

-5x-2=13 we can add 2 to both sides of the equation reason y... -2+2=0

OpenStudy (lynfran):

so we have -5x=13+2 whats next??

OpenStudy (anonymous):

13+2 equals 15?

OpenStudy (lynfran):

yes -5x=15 now how can we get the x term alone ...since -5 is being multiple to x giving -5x

OpenStudy (lynfran):

whats the opposite of multiplication?

OpenStudy (anonymous):

division

OpenStudy (lynfran):

yes we divide both sides of the equation by.... ??

OpenStudy (anonymous):

5?

OpenStudy (lynfran):

shd be -5

OpenStudy (lynfran):

if we divide by 5 then we have -x=15/5

OpenStudy (anonymous):

so x equals -3

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

Could you help me with -1=4+x/3?

OpenStudy (lynfran):

what the first step...do u know..?

OpenStudy (anonymous):

subtract 4 from each side

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

so, x/3=-5

OpenStudy (lynfran):

yes now whats the next step?

OpenStudy (anonymous):

multiply 3 to each side

OpenStudy (lynfran):

correct

OpenStudy (anonymous):

so, x=-15?

OpenStudy (lynfran):

yep

OpenStudy (anonymous):

Thank you very much for helping me!

OpenStudy (lynfran):

welcome